# Herding Unmasked: Insights into Cryptocurrencies, Stocks and US ETFs

### v1

@@ 5 Conclusions
This work aims to detect and analyze investor herding behavior in different types of investment vehicle separately and also in their combination, including stocks, US ETFs and cryptocurrencies. We enrich the existing literature with three contributions: Firstly, the use of a more up-to-date and granular historical price time series covering several recent events, i.e. the US-China trade war in 2019, the Covid-19 pandemic followed by the global economic crisis in 2020, the Bull market period in 2021-2022 and the Ukraine-Russia conflict which started in 2022. Secondly, the use of multiple investment vehicle types adds to the absence of diversified market issues in the existing studies. Thirdly, the focus on herding behavior at a community level (small scale) enhances the clarity and granularity of this herding phenomenon.
@@
Regarding the first research question, when examining the herding behavior in each type of investment vehicle separately, we find a similar herding pattern between stocks and US ETFs, with evidence of herding during the Pre-Covid-19 coinciding with the US-China trade war and during the Covid-19 pandemic. By contrast, the cryptocurrency market displays quite distinct behavior since its investors are seen to have performed herding during the Covid-19 pandemic, the Post-Covid-19 when the market was recovering and also during the Bull market period when assets experienced a surge in prices. In addition, the magnitude of herding in the US ETFs is much stronger than that in stocks and cryptocurrencies.
@@
Regarding the second research question, we execute a deeper herding detection and analysis by breaking down the whole combination of stocks, cryptocurrencies and US ETFs) into smaller communities where each of these comprises the assets that are most similar to each other. Additionally, for each community, we also use information about sector distributions, with each asset belonging to one of the 13 sectors, including cryptocurrency, US ETF and 11 stock sectors. Our main finding is that herding behavior exists at all times but at a smaller scale (i.e. community level), in all types of investment vehicle examined. This mainly comes from specific events which particularly impact on that community. As a result, although herding might not be seen in a whole market composed of all assets, it can be found by looking at a particular community/subset of assets. Moreover, we notice 4 stock sectors in which for each sector, its assets and relevant ETFs tend to form communities among themselves, with the herding behavior found in these communities driven by unique events only influencing that sector. These include Energy, Healthcare, Information Technology and Finance. On the other hand, this characteristic does not appear in the remaining sectors, i.e. Industry, Consumer Discretionary and Consumer Staples.
@@
This work enriches the understanding of herding tendencies in the examined types of investment vehicle that help investors construct an appropriate investment strategy. Furthermore, we also raise investors’ attention towards 4 stock sectors, i.e. Energy, Healthcare, Information Technology and Finance as these seem more sensitive to herding behavior than other sectors such as Industry, Consumer Discretionary and Consumer Staples. Therefore, investors who invest in these former four stock sectors should be more cautious with relevant economic and political turmoil since they are more likely to form a herding behavior during such occasions. Future work could expand on this study by using a larger dataset, covering more types of investment vehicle, such as commodities, currencies and bonds. Another area of future research could involve different time granularities. Furthermore, several findings of this study warrant further investigation: Firstly, the differences in herding magnitude between different communities. Secondly, the comparison between CSAD, CSAD for up and CSAD for down markets within each community, between different communities and between different sub-periods. Thirdly, the connections between ETFs tracking the S&\&&P-related indices and technology stocks.
Acknowledgments
The author An Pham Ngoc Nguyen wishes to acknowledge the financial support of the Science Foundation Ireland Centre for Research Training in Artificial Intelligence under grant number 18/CRT/6223. The authors Martin Crane and Marija Bezbradica wish to acknowledge the support from the Science Foundation Ireland under Grant Agreement No. 13/RC/2106_P2 at the ADAPT SFI Research Centre at DCU. ADAPT, the SFI Research Centre for AI-Driven Digital Content Technology, is funded by the Science Foundation Ireland through the SFI Research Centres Programme.
Appendix
Appendix A Figure Clarification
Clarification 1.
The process of community detection. Firstly, a set of time series is determined, where each time series corresponds to an asset. Secondly, a Pearson-based correlation matrix is constructed from the historical time series. the darker the color, the more similar the two corresponding assets. Thirdly, a distance matrix is obtained from the correlation matrix. Fourthly, a fully connected graph is constructed based on the distance between each pair of assets from the previous step. Specifically, although every two arbitrary nodes are connected, their corresponding edge has a different weight which represents the distance between them, the lower the weight, the nearer the two nodes. Fifthly, a minimum spanning tree (MST) is extracted from the distance matrix to reduce the size of the graph while keeping essential information. Lastly, Louvain community detection technique is applied to the MST to detect existing communities in the graph. Each community contains nodes that share similar characteristics and nodes in different communities have different characteristics.
Clarification 2.
Visualization of herding detection results and sector distribution. The figure illustrates herding detection results and sector distribution of each community. a) herding results for each community. The grey bar means the herding based on CSAD model (Eq 3) is significant, the green bar means the herding based on CSAD for the up market condition (Eq 4) is significant, the red bar means the herding based on CSAD for down market condition (Eq 4) is significant. The higher the bar, the more pronounced the herding; b) sector distribution (percentage of each sector) of each community, sectors are listed on the y-axis, each represented by a unique color. For each community, the longer the bar of a corresponding sector, the greater the sector’s percentage within that community.
Clarification 3.
Time-varying community structures of cryptocurrencies, stocks and US ETFs. In each sub-figure, assets with the same color belong to one community. The black-shaped oval highlights an example of a community containing stocks and US ETFs where the centre of the community is a US ETF, which is linked by surrounding stocks. Moreover, this US ETF tends to be the basket of its surrounding stocks, meaning that it is designed to track the overall performance of those stocks. The purple-shaped oval highlights communities containing cryptocurrencies. As can be seen, most cryptocurrencies are close to each other, forming communities among themselves and showing a certain distinction level from stocks and US ETFs.
Appendix B Investment Vehicles Used in this Study and Their Corresponding Assets
Table 6 introduces three types of investment vehicle used in our study and displays available assets in each type of investment vehicle.
Table 6: Three types of investment vehicle used in this study and their corresponding assets
Type Assets
Cryptocurrencies ADA (Cardano), BAT (Basic Attention Token), BCH (Bitcoin Cash), BNT (Bancor), BSV (Bitcoin SV), BTC (Bitcoin), DAI (Dai), DASH (Dash), DOGE (Dogecoin), EOS (EOS), ETH (Ethereum), ETC (Ethereum Classic), LTC (Litecoin), NEO (Neo), OMG (OMG Network), QTUM (Qtum), REP (Augur), TRX (TRON), USDT (Tether), UST (TerraUSD), XLM (Stellar), XMR (Monero), XRP (Ripple), XTZ (Tezos), XVG (Verge), ZEC (Zcash), ZRX (Ox)
US ETFs CIBR (First Trust Nasdaq Cybersecurity ETF), DIA (SPDR Dow Jones Industrial Average ETF), DJD (Invesco Dow Jones Industrial Average Dividend ETF), DJP (iPath Bloomberg Commodity Index Total Return ETN), DVY (iShares Select Dividend ETF), FDLO (Fidelity Low Volatility Factor ETF), FDN (First Trust Dow Jones Internet Index Fund), FNGU (MicroSectors FANG+ Index 3X Leveraged ETN), FTXO (First Trust Nasdaq Bank ETF), IBB (iShares Nasdaq Biotechnology ETF), IDU (iShares U.S. Utilities ETF), IEO (iShares U.S. Oil & Gas Exploration & Production ETF), IJH (iShares Core S&P Mid-Cap ETF), IJJ (iShares S&P Mid-Cap 400 Value ETF), IJK (iShares S&P Mid-Cap 400 Growth ETF), IJR (iShares Core S&P Small-Cap ETF), IJS (iShares S&P Small-Cap 600 Value ETF), IJT (iShares S&P Small-Cap 600 Growth ETF), ITB (iShares U.S. Home Construction ETF), ITOT (iShares Core S&P Total U.S. Stock Market ETF), IUSG (iShares Core S&P U.S. Growth ETF), IVE (iShares S&P 500 Value ETF), IWB (iShares Russell 1000 ETF), IWM (iShares Russell 2000 ETF), IWV (iShares Russell 3000 ETF), IYE (iShares U.S. Energy ETF), IYF (iShares U.S. Financials ETF), IYJ (iShares U.S. Industrials ETF), IYR (iShares U.S. Real Estate ETF), IYT (iShares Transportation Average ETF), IYW (iShares U.S. Technology ETF), IYY (iShares Dow Jones U.S. ETF), KBWB (Invesco KBW Bank ETF), OEF (iShares S&P 100 ETF), ONEQ (Fidelity Nasdaq Composite ETF), PNQI (Invesco Nasdaq Internet ETF), QQQ (Invesco QQQ Trust), QTEC (First Trust NASDAQ-100 Technology Sector Index Fund), QYLD (Global X NASDAQ 100 Covered Call ETF), ROBT (First Trust Nasdaq Artificial Intelligence and Robotics ETF), RPG (Invesco S&P 500 Pure Growth ETF), RPV (Invesco S&P 500 Pure Value ETF), SLYG (SPDR S&P 600 Small Cap Growth ETF), SLYV (SPDR S&P 600 Small Cap Value ETF), SLY (SPDR S&P 600 Small Cap ETF), SOXX (iShares PHLX Semiconductor ETF), SPY (SPDR S&P 500 ETF), SVXY (ProShares Short VIX Short-Term Futures ETF), XTN (SPDR S&P Transportation ETF).
Stocks AAPL (Apple Inc.), ABBV (AbbVie Inc.), ABB (ABB Ltd.), ABT (Abbott Laboratories), ACN (Accenture plc), ADBE (Adobe Inc.), ADI (Analog Devices, Inc.), ADP (Automatic Data Processing, Inc.), AIR (AAR Corp.), ALV (Autoliv Inc.), AMAT (Applied Materials, Inc.), AMD (Advanced Micro Devices, Inc.), AMGN (Amgen Inc.), AMT (American Tower Corporation), AMZN (Amazon.com Inc.), ASML (ASML Holding NV), ASX (ASE Technology Holding Co., Ltd.), AVGO (Broadcom Inc.), AXP (American Express Company), AZN (AstraZeneca PLC), BABA (Alibaba Group Holding Limited), BAC (Bank of America Corp.), BA (The Boeing Company), BDX (Becton, Dickinson and Company), BHP (BHP Group Limited), BKNG (Booking Holdings Inc.), BKR (Baker Hughes Company), BLK (BlackRock Inc.), BMY (Bristol Myers Squibb Company), BP (BP plc), BSX (Boston Scientific Corporation), BTI (British American Tobacco plc), BUD (Anheuser-Busch InBev SA/NV), BX (The Blackstone Group Inc.), CAT (Caterpillar Inc.), CB (Chubb Limited), CFR (Cullen/Frost Bankers, Inc.), CI (Cigna Corporation), CMCSA (Comcast Corporation), CNI (Canadian National Railway Company), COP (ConocoPhillips), COST (Costco Wholesale Corporation), CP (Canadian Pacific Railway Limited), CRM (Salesforce.com Inc.), CSCO (Cisco Systems, Inc.), CSL (Carlisle Companies Incorporated), CVS (CVS Health Corporation), CVX (Chevron Corporation), C (Citigroup Inc.), DEO (Diageo plc), DE (Deere & Company), DHR (Danaher Corporation), DIS (The Walt Disney Company), DTE (DTE Energy Company), ELV (Eaton Vance Corporation), EL (The Estée Lauder Companies Inc.), EOG (EOG Resources Inc.), EQNR (Equinor ASA), ETN (Eaton Corporation plc), GE (General Electric Company), GILD (Gilead Sciences, Inc.), GOOG (Alphabet Inc.), GS (The Goldman Sachs Group, Inc.), HCA (HCA Healthcare, Inc.), HDB (HDFC Bank Limited), HD (The Home Depot, Inc.), HON (Honeywell International Inc.), HSBC (HSBC Holdings plc), IBM (International Business Machines Corporation), IBN (ICICI Bank Limited), IDEX (Ideanomics, Inc.), INTC (Intel Corporation), INTU (Intuit Inc.), ISRG (Intuitive Surgical, Inc.), ITW (Illinois Tool Works Inc.), JNJ (Johnson & Johnson), JPM (JPMorgan Chase & Co.), KO (The Coca-Cola Company), LIN (Linde plc), LLY (Eli Lilly and Company), LMT (Lockheed Martin Corporation), LOW (Lowe’s Companies, Inc.), LRCX (Lam Research Corporation), MA (Mastercard Incorporated), MCD (McDonald’s Corporation), MC (Moelis & Company), MDLZ (Mondelez International, Inc.), MDT (Medtronic plc), META (Meta Platforms, Inc.), MMC (Marsh & McLennan Companies, Inc.), MO (Altria Group, Inc.), MRK (Merck & Co., Inc.), MSFT (Microsoft Corporation), MS (Morgan Stanley), MUFG (Mitsubishi UFJ Financial Group, Inc.), MU (Micron Technology, Inc.), NEE (NextEra Energy, Inc.), NFLX (Netflix, Inc.), NKE (NIKE, Inc.), NOW (ServiceNow, Inc.), NVDA (NVIDIA Corporation), NVO (Novo Nordisk A/S), NVR (NVR, Inc.), NVS (Novartis AG), ORCL (Oracle Corporation), PANW (Palo Alto Networks, Inc.), PBR (Petroleo Brasileiro S.A. - Petrobras), PDD (Pinduoduo Inc.), PEP (PepsiCo, Inc.), PFE (Pfizer Inc.), PG (Procter & Gamble Co.), PLD (Prologis, Inc.), PM (Philip Morris International Inc.), PYPL (PayPal Holdings, Inc.), QCOM (QUALCOMM Incorporated), SAP (SAP SE), SBUX (Starbucks Corporation), SCHW (The Charles Schwab Corporation), SHEL (Sealed Air Corporation), SHOP (Shopify Inc.), SLB (Schlumberger Limited), SNY (Sanofi), SONY (Sony Group Corporation), SO (The Southern Company), SYK (Stryker Corporation), TD (The Toronto-Dominion Bank), TJX (The TJX Companies, Inc.), TMO (Thermo Fisher Scientific Inc.), TMUS (T-Mobile US, Inc.), TM (Toyota Motor Corporation), TSLA (Tesla, Inc.), TSM (Taiwan Semiconductor Manufacturing Company Limited), TTE (TotalEnergies SE), TXN (Texas Instruments Incorporated), T (AT&T Inc.), UL (Unilever PLC), UNH (UnitedHealth Group Incorporated), UNP (Union Pacific Corporation), UPS (United Parcel Service, Inc.), VRTX (Vertex Pharmaceuticals Incorporated), VZ (Verizon Communications Inc.), V (Visa Inc.), WFC (Wells Fargo & Co.), WMT (Walmart Inc.), XOM (Exxon Mobil Corporation), ZTS (Zoetis Inc.)
Appendix C Classification of Financial Sectors in the Stock Market
Table 7 lists all stock sectors in the stock market and available assets from our dataset in each sector.
Table 7: Distribution of Stock Sectors in the Stock Dataset. Each Asset is Assigned to an Appropriate Stock Sector
Information
Technology
Healthcare Financials
Consumer
discretionary
Consumer
staples
Industrials Energy Utilities Materials Real estate
Communication
services
ASX
AAPL
ACN
ADBE
ADI
ADP
AMAT
AMD
ASML
AVGO
BABA
CMCSA
CRM
CSCO
GOOG
IBM
INTC
INTU
LRCX
MA
META
MSFT
MU
NFLX
NOW
NVDA
ORCL
PANW
PDD
QCOM
SAP
SHOP
SONY
TMUS
TSM
TXN
T
VZ
V
ABBV
ABT
AMGN
AZN
BDX
BMY
BSX
CI
CVS
DHR
ELV
GILD
HCA
ISRG
JNJ
LLY
MDT
MRK
NVO
NVS
PFE
SNY
SYK
TMO
UNH
VRTX
ZTS
AXP
BAC
BLK
BX
CB
CFR
C
GS
HDB
HSBC
IBN
JPM
MC
MMC
MS
MUFG
SCHW
TD
WFC
ALV
AZMN
BKNG
COST
CSL
DIS
HD
IDEX
LOW
MCD
NKE
NVR
SBUX
TJX
TM
TSLA
BTI
BUD
DEO
EL
GE
HON
ITW
KO
MDLZ
MO
PEP
PG
PM
UL
WMT
ABB
AIR
BA
CAT
CNI
CP
DE
ETN
LMT
UNP
UPS
PYPL
BKR
BP
COP
CVX
EQNR
PBR
SHEL
SLB
TTE
EOG
XOM
DTE
NEE
SO
BHP
LIN
AMT
PLD
Appendix D Details of Assets in each Community (Entire Period)
Table 8 provides the result of community detection on the combination of stocks, cryptocurrencies and US ETFs throughout the entire period from 04/2019 until 05/2023.
Table 8: Assets in each community over the entire period. The red color indicates cryptocurrencies, the blue color indicates US ETFs and the black color indicates stocks
Community Assets
1 AAPL, ADBE, AMGN, GILD, GOOG, INTU, MSFT, SONY, TSLA, VRTX, FNGU, IBB, IYW, ONEQ, QQQ, QYLD
2 ABBV, AMT, BMY, BSX, CMCSA, ETN, HCA, HON, IBM, ITW, KO, LLY, MDLZ, MDT, MMC, PEP, PG, PLD, SYK, DJD, IVE, IYJ, IYR
3 ABB, ACN, BLK, BX, DIS, EL, ISRG, MA, NKE, ORCL, SAP, SBUX, TMUS, V, ZTS, FDLO, ITOT, IWB, IWV, IYY, SPY, SVXY
4 BTI, CAT, CB, DE, DTE, LMT, MO, NEE, PM, SO, T, VZ, DVY, IDU
5 ADI, AMAT, AMD, ASML, ASX, AVGO, INTC, LRCX, MU, NVDA, QCOM, TSM, TXN, QTEC, ROBT, SOXX
6 ALV, BHP, BKNG, CNI, CP, CSL, MC, NVR, TJX, UNP, UPS, IJH, IJJ, IJK, ITB, IYT, XTN
7 ABT, BDX, COST, DHR, TMO, WMT, BNT, BTC, DOGE, ETH, USDT, XVG, IUSG, OEF, RPG
8 AIR, IJR, IJS, IJT, IWM, SLYG, SLYV, SLY
9 AMZN, BABA, CRM, IDEX, META, NFLX, NOW, PANW, PDD, PYPL, SHOP, CIBR, FDN, PNQI
10 ADA, BAT, BCH, BSV, DAI, DASH, EOS, ETC, LTC, NEO, OMG, QTUM, REP, TRX, XLM, XMR, XRP, XTZ, ZEC, ZRX
11 ADP, AZN, BA, BUD, CI, CSCO, CVS, DEO, ELV, HD, JNJ, LIN, LOW, MCD, MRK, NVO, NVS, PFE, SNY, TM, UL, UNH, DIA
12 AXP, BAC, CFR, C, GS, HDB, HSBC, IBN, JPM, MS, MUFG, SCHW, TD, WFC, FTXO, IYF, KBWB
13 BKR, BP, COP, CVX, EOG, EQNR, GE, PBR, SHEL, SLB, TTE, XOM, DJP, IEO, IYE, RPV
Appendix E Details of Assets in each Community (Sub-Period)
Tables 9 - 13 provide the results of community detection on the combination of stocks, cryptocurrencies and US ETFs during different sub-periods, including Pre-Covid-19 (04/2019 - 12/2019), Covid-19 pandemic (01/2020 - 06/2020), Post-Covid-19 (07/2020 - 12/2020), Bull market time (01/2021 - 04/2022), Ukraine-Russia conflict (05/2022 - 05/2023).
Table 9: Assets in each community during the Pre-Covid-19. The red color indicates cryptocurrencies, the blue color indicates US ETFs and the black color indicates stocks
Community Assets
1 ACN, ADP, AMT, DTE, EL, INTU, ISRG, LMT, MA, MCD, NEE, NOW, PLD, PYPL, SAP, SBUX, SONY, SO, V, IDU, IUSG, IYR
2 ABBV, AXP, BUD, CMCSA, COST, HON, LIN, MMC, TJX, WMT, OEF, SPY, SVXY
3 ABB, BHP, BLK, CAT, CB, CI, CVS, DE, ELV, ETN, GE, HSBC, ITW, MUFG, TD, UNH, IVE, IYF
4 ABT, BDX, BSX, DHR, HDB, IBN, LLY, MDT, MRK, NVS, PFE, SYK, TMO, ZTS
5 BAT, BNT, DAI, DASH, ETH, OMG, USDT, UST, XLM, XMR, XRP, XVG, ZEC, ZRX
6 ADBE, CRM, GOOG, META, PANW, SHOP, CIBR, FDN
7 ADI, AMAT, AMD, ASX, AVGO, INTC, LRCX, MU, NVDA, QCOM, TSM, TXN, SOXX
8 AIR, BX, CNI, CP, CSL, MC, NVR, ORCL, TMUS, UNP, UPS, IJH, IJJ, IJK, ITB, ITOT, IYT, XTN
9 ALV, BAC, CFR, C, GS, JPM, MS, SCHW, WFC, FTXO, KBWB, RPV
10 AMGN, BMY, GILD, HCA, VRTX, IBB, IJR, IJS, IJT, IWM, SLYG, SLYV, SLY
11 AZN, DEO, KO, MDLZ, NVO, PEP, PG, SNY, TM, UL, FDLO, IWB, IWV, IYJ, IYY, RPG
12 ADA, BCH, BSV, BTC, DOGE, EOS, ETC, LTC, REP
13 AAPL, AMZN, ASML, BABA, BKNG, CSCO, IDEX, MSFT, NFLX, PDD, TSLA, FNGU, IYW, ONEQ, PNQI, QQQ, QTEC, QYLD, ROBT
14 BA, DIS, HD, IBM, JNJ, LOW, NKE, DIA, DJD
15 BKR, BP, COP, CVX, EOG, EQNR, PBR, SHEL, SLB, TTE, XOM, DJP, IEO, IYE
16 BTI, MO, PM, T, VZ, NEO, QTUM, TRX, XTZ, DVY
Table 10: Assets in each community during the Covid-19 pandemic. The red color indicates cryptocurrencies, the blue color indicates US ETFs and the black color indicates stocks
Community Assets
1 AAPL, ADBE, AMGN, AMZN, COST, GILD, GOOG, INTC, MSFT, PYPL, SONY, TMUS, VRTX, WMT, IBB, QQQ, QYLD
2 BTI, CB, DE, DTE, ETN, MO, NEE, SO, T, VZ, DVY, IDU
3 ABB, BUD, CI, DEO, DIS, ELV, EL, HD, LOW, MCD, NKE, SBUX, UNH, DIA
4 ABBV, ABT, AZN, BDX, CSCO, DHR, IDEX, JNJ, MA, MRK, NVO, NVS, ORCL, PFE, SNY, TMO, V, IUSG, OEF, ONEQ, SVXY
5 ADI, AMAT, AMD, ASML, ASX, AVGO, LRCX, MU, NVDA, QCOM, TSM, TXN, SOXX
6 ADP, BSX, CAT, CMCSA, CNI, CP, CVS, GE, HON, IBM, ITW, KO, LMT, PM, UNP, UPS, DJD, IVE, IYJ, IYT, XTN
7 INTU, NOW, ADA, BAT, BNT, ETH, NEO, OMG, QTUM, XMR, XTZ, XVG, ZRX, IYW, QTEC, RPG
8 ALV, BA, BHP, BKR, BP, COP, CVX, EOG, EQNR, MUFG, PBR, SHEL, SLB, TD, TTE, XOM, DJP, IEO, IYE, RPV
9 AMT, AXP, BAC, BX, CFR, C, GS, HDB, HSBC, IBN, JPM, MS, PLD, SCHW, TJX, WFC, FTXO, IYF, IYR, KBWB
10 BABA, CRM, META, NFLX, PANW, PDD, SHOP, TSLA, CIBR, FDN, FNGU, PNQI
11 BCH, BSV, BTC, DAI, DASH, DOGE, EOS, ETC, LTC, REP, TRX, XLM, XRP, ZEC
12 AIR, BKNG, CSL, HCA, MC, NVR, TM, IJH, IJJ, IJK, IJR, IJS, IJT, ITB, IWM, ROBT, SLYG, SLYV, SLY
13 ACN, BLK, BMY, ISRG, LIN, LLY, MDLZ, MDT, MMC, PEP, PG, SAP, SYK, UL, ZTS, USDT, UST, FDLO, ITOT, IWB, IWV, IYY, SPY
Table 11: Assets in each community during the Post-Covid-19. The red color indicates cryptocurrencies, the blue color indicates US ETFs and the black color indicates stocks
Community Assets
1 BX, ISRG, FDLO, ITOT, IUSG, IWB, IWV, IYY, OEF, ROBT, RPG
2 ABBV, AZN, BMY, GILD, JNJ, LLY, MRK, NVO, NVS, PFE, SNY
3 ABB, ABT, ACN, BDX, BLK, DHR, HD, LOW, MA, ORCL, TMO, V, ZTS, SPY, SVXY
4 ADBE, CRM, META, NFLX, NOW, PANW, PDD, PYPL, SHOP, CIBR, FDN, PNQI
5 ADI, AMAT, ASML, ASX, AVGO, INTC, LRCX, MU, QCOM, TSM, TXN, SOXX
6 ADP, BSX, CAT, CMCSA, CSCO, CVS, DE, ETN, HON, IBM, ITW, KO, LMT, MDLZ, MDT, MMC, PEP, PG, SYK, TJX, DJD, IVE
7 AIR, BA, CSL, GE, HCA, MC, MUFG, TM, IJJ, IJS, RPV, SLYV
8 CI, DIS, ELV, EL, HDB, IBN, LIN, MCD, NKE, SAP, SBUX, UNH, DIA
9 ALV, CNI, CP, NVR, UNP, UPS, IJH, IJK, IJR, IJT, ITB, IWM, IYJ, IYT, SLYG, SLY, XTN
10 AMGN, AMZN, BABA, COST, GOOG, INTU, SONY, TMUS, TSLA, VRTX, WMT, FNGU, IBB, ONEQ, QQQ, QYLD
11 AMT, AXP, BKNG, GS, MS, PLD, IYF, IYR
12 BAC, CFR, C, JPM, SCHW, WFC, DAI, FTXO, KBWB
13 IDEX, BAT, BCH, BSV, DOGE, EOS, ETC, NEO, OMG, QTUM, REP, TRX, XLM, XRP, XTZ, XVG, ZRX
14 AAPL, AMD, MSFT, NVDA, ADA, BNT, BTC, DASH, ETH, LTC, XMR, ZEC, IYW, QTEC
15 BHP, BKR, BP, BUD, COP, CVX, DEO, EOG, EQNR, PBR, SHEL, SLB, TTE, UL, XOM, DJP, IEO, IYE
16 BTI, CB, DTE, HSBC, MO, NEE, PM, SO, TD, T, VZ, USDT, UST, DVY, IDU
Table 12: Assets in each community during the Bull market time. The red color indicates cryptocurrencies, the blue color indicates US ETFs and the black color indicates stocks
Community Assets
1 AAPL, ADBE, AMZN, GOOG, INTU, MSFT, TSLA, FNGU, IYW, ONEQ, QQQ, QTEC, QYLD
2 AMT, BTI, CMCSA, CVS, DTE, ETN, HCA, ITW, KO, MO, NEE, PLD, PM, SO, DAI, DJD, IDU, IVE, IYR
3 CNI, CP, MA, UNP, UPS, V, IYJ, IYT, XTN
4 CAT, DE, GE, IBM, LMT, T, VZ, DVY, IJJ, RPV
5 ADI, AMAT, AMD, ASML, ASX, AVGO, INTC, LRCX, MU, NVDA, QCOM, TSM, TXN, SOXX
6 AIR, ALV, BA, CSL, HD, IDEX, LOW, MC, NVR, IJH, IJK, IJR, IJS, IJT, ITB, IWM, SLYG, SLYV, SLY
7 ABT, BDX, DHR, ISRG, SAP, TMO, VRTX, ZTS, IBB, IUSG, OEF, ROBT, RPG
8 ABBV, AMGN, AZN, BMY, GILD, JNJ, LLY, MDLZ, MRK, NVO, NVS, PEP, PFE, PG, SNY, UL, UST
9 ABB, ACN, ADP, BLK, BSX, COST, EL, MDT, ORCL, SYK, TMUS, WMT, FDLO,SPY, SVXY
10 AXP, BAC, CB, CFR, C, GS, HSBC, JPM, MS, MUFG, SCHW, WFC, FTXO, IYF, KBWB
11 BABA, CRM, META, NFLX, NOW, PANW, PDD, PYPL, SHOP, CIBR, FDN, PNQI
12 BHP, BKR, BP, COP, CVX, EOG, EQNR, PBR, SHEL, SLB, TTE, XOM, DJP, IEO, IYE
13 BKNG, BUD, CI, CSCO, DEO, ELV, HON, LIN, MCD, MMC, SBUX, TJX, TM, UNH, DIA
14 TD, ADA, BAT, BCH, BSV, DASH, EOS, ETC, NEO, OMG, QTUM, REP, TRX, USDT, XLM, XMR, XRP, XTZ, ZEC
15 BX, DIS, HDB, IBN, NKE, SONY, ITOT, IWB, IWV, IYY
16 BNT, BTC, DOGE, ETH, LTC, XVG, ZRX
Table 13: Assets in each community during the Ukraine-Russia Conflict. The red color indicates cryptocurrencies, the blue color indicates US ETFs and the black color indicates stocks
Community Assets
1 BHP, BKR, BP, COP, CVX, EOG, EQNR, PBR, SHEL, SLB, TTE, XOM, DJP, IEO, IYE
2 ABBV, BMY, JNJ, KO, LLY, MDLZ, MO, MRK, PEP, PFE, PG, PM
3 AMT, BTI, CAT, CB, CMCSA, DE, DTE, ITW, LMT, MMC, NEE, PLD, SO, T, VZ, DVY, IDU, IVE, IYR, RPV
4 ABT, ACN, BDX, BLK, BSX, BX, DIS, EL, HDB, IBN, ISRG, MA, MDT, NKE, SAP, SBUX, SONY, SYK, V, ITOT, SPY, SVXY
5 ADBE, ADI, AMAT, AMD, ASML, ASX, AVGO, GOOG, IDEX, INTC, LRCX, MSFT, MU, NVDA, QCOM, TSM, TXN, IYW, QTEC, ROBT, SOXX
6 BUD, CI, COST, CSCO, CVS, DEO, ELV, HON, IBM, LIN, MCD, TMUS, UL, UNH, WMT, DIA, DJD
7 ADP, ORCL, TM, ZTS, FDLO, IWB, IWV, IYJ, IYY, RPG
8 AIR, IJR, IJS, IJT, IWM, SLYG, SLYV, SLY
9 AMGN, AZN, GILD, NVO, NVS, SNY, VRTX, IBB
10 ABB, ALV, BA, BKNG, CNI, CP, CSL, ETN, GE, HCA, HD, LOW, MC, NVR, TJX, UNP, UPS, IJH, IJJ, IJK, ITB, IYT, XTN
11 AMZN, CRM, NFLX, NOW, PANW, PYPL, SHOP, CIBR, FDN, PNQI
12 AXP, BAC, CFR, C, GS, HSBC, JPM, MS, MUFG, SCHW, TD, WFC, DAI, FTXO, IYF, KBWB
13 ADA, BCH, BNT, BSV, BTC, DASH, DOGE, EOS, ETC, ETH, LTC, NEO, OMG, QTUM, TRX, USDT, XLM, XMR, XRP, XVG, ZEC
14 AAPL, BABA, DHR, INTU, META, PDD, TMO, TSLA, BAT, REP, XTZ, ZRX, FNGU, IUSG, OEF, ONEQ, QQQ, QYLD

### v2

@@ 5 Conclusions
This work aims to detect and analyze investor herding behavior in different types of investment vehicles separately and also in their combination, including stocks, US ETFs and cryptocurrencies. We enrich the existing literature with three contributions: Firstly, the use of a more up-to-date and granular historical price time series covering several recent events, i.e. the US-China trade war in 2019, the Covid-19 pandemic followed by the global economic crisis in 2020, the Bull market period 2021-2022 and the Ukraine-Russia conflict which started in 2022, provides the latest insights into herding behavior in financial markets at a finer scale (30-min granularity). Secondly, the use of multiple investment vehicle types, enables us to examine herding behavior within a wider market context. Thirdly, the focus on herding behavior at a community level, enhances the clarity and granularity of this herding phenomenon.
@@
Regarding the first research question, when examining the herding behavior in each type of investment vehicle separately, we find similar herding patterns between stocks and US ETFs for the majority of the time considered, with evidence of herding in both during the Pre-Covid-19 coinciding with the US-China trade war and during the Covid-19 pandemic. By contrast, the cryptocurrency market displays quite distinct behavior since its investors are seen to have herded during the Covid-19 pandemic, the first Bull Time when the market was recovering, the second Bull Time when most assets experienced a surge in prices and also the second Ukraine-Russia Conflict when another upward trend in financial markets took place.
@@
Regarding the second research question, we execute a deeper herding detection and analysis by breaking down the whole combination of stocks, cryptocurrencies and US ETFs) into smaller communities where each of these comprises the assets that are most similar to each other. Additionally, for each community, we also use information about sector distributions, with each asset belonging to one of the 13 sectors, including cryptocurrency, US ETF and 11 stock sectors. Our main finding is that herding behavior exists at all times but at a smaller scale (i.e. community level), in all types of investment vehicle examined. This mainly comes from specific events which particularly impact on that community. As a result, although herding might not be seen in a whole market composed of all assets, it can be found by looking at a particular community/subset of assets. Moreover, we notice 4 stock sectors in which for each sector, its assets and relevant ETFs tend to form communities among themselves, with the herding behavior found in these communities driven by unique events only influencing that sector. These include Energy, Healthcare, Information Technology and Financials. On the other hand, this characteristic does not appear in the remaining sectors, i.e. Industry, Consumer Discretionary and Consumer Staples.
@@
This work considerably adds to the body of knowledge around understanding herding tendencies in the examined types of investment vehicle, principally to help investors construct an appropriate investment strategy. Apart from several factors causing herding behavior within a market such as investor sentiment, macroeconomics, major impactful events, news, and internal price movements, US ETFs seem to be another herding driver in financial markets, especially in the direction of cryptocurrencies. This is reflected in the significant contagion effects that US ETFs exert on both stocks and cryptocurrencies. Additionally, contagion from stocks to US ETFs is consistently stronger than from cryptocurrencies to US ETFs. This tighter connection between US ETFs and stocks explains why they tend to form communities together. Furthermore, we also raise investors’ attention towards 4 stock sectors, i.e. Energy, Healthcare, Information Technology and Financials as these seem more sensitive to herding behavior than other sectors such as Industry, Consumer Discretionary and Consumer Staples. Therefore, investors who invest in these former four stock sectors should be more cautious with relevant economic and political turmoil since they are more likely to form a herding behavior during such occasions. Future work could expand on this study by using a larger dataset, covering more types of investment vehicle, such as commodities, currencies and bonds. Another area of future research could involve different time granularities. Furthermore, several findings of this study warrant further investigation: Firstly, the differences in herding magnitude between different communities. Secondly, the comparison between CSAD, CSAD for up and CSAD for down markets within each community, between different communities and between different sub-periods. Thirdly, the connections between ETFs tracking the S&\&&P-related indices and technology stocks.
Supporting Information
S1 Appendix.
The process of community detection.
S1 Text.
Visualization of herding detection results and sector distribution.
S2 Text.
Time-varying community structures of cryptocurrencies, stocks and US ETFs.
S1 Table.
List of Assets Used in this Study.
S2 Table.
Classification of Stock Sectors in the Stock Markets.
S3 Table.
Details of Assets in each Community (Entire Period).
S4 Table.
Details of Assets in each Community (Sub-Period).
S1 File.
Contagion Effects between Cryptocurrencies, Stocks and US ETFs in each Sub-period.